# Split the Yield, Share the Risk: Pricing, Hedging and Fixed rates in DeFi

### v2

@@ Abstract
keywords:
Decentralized Finance, Lending, Automated Market Makers, Fixed Rates, Yield Tokenization, Liquid Staking
@@ 1 Introduction
Although we focus on DeFi lending as the primary case study of where yield tokenization can be used, our results can be extended to other yield-bearing actions such as AMMs and staking as well (Appendix [H]). In automated market makers, liquidity providers (LPs) contribute assets to a pool, earning fees from trades that constitute the yield on their LP tokens. By tokenizing this yield, LPs can sell it to secure a steady cash flow. A more powerful strategy is hedging impermanent loss (loss-vs-holding) [49], by connecting yield tokenization to a demand lending pool [49, 17], allowing LPs to borrow against their LP positions and shift the risk of asset price fluctuation to the lending pool manager. Alternatively, the liquid markets of yield tokens based on LP positions may enable estimation of arbitrage loss (loss-vs-rebalancing) [57], and hedge against uncertainty in the earning of swap fees to compensate for this loss [57]. In liquid staking, tokenizing staking yield similarly allows risk management, though it also introduces some negative externalities (Appendix [H]). A liquid market for such yield tokens enables adversarial validators to profit by shorting the yield and misbehaving in consensus protocols, exacerbating the principal-agent problem that is present there [21]. Nonetheless, yield tokenization enables hedging against other risks, such as Ethereum’s congestion fees: since variable staking rewards depend on transaction/blob fees and validators are known in advance, buyers of yield tokens can hedge against congestion costs while delegators use them to lock in staking rewards [40]. This dual-sided risk transfer resembles the borrower-lender dynamics that we analyze in Section 4 and can be further improved with AMM designs tailored to participants’ risk preferences.
@@
In this work, we make the following key contributions. Firstly, we propose a mathematical model of yield tokenization using stochastic differential equations (Section 2.1). To our knowledge, this is the first formal treatment of yield tokenization and can be used to characterize any yield-bearing token. Secondly, given this model, we derive the fair price of a yield token from a trader’s perspective, in a sense that ensures no arbitrage (Section 2.2). Thirdly, we show how lenders and borrowers in DeFi lending protocols can use yield tokenization to hedge against interest rate uncertainty - thereby improving the welfare of all lending pool participants (Section 3). Fourthly, we design a market maker (using a menu of bonding curves) for yield tokens to be traded efficiently, so that the hedging that we mentioned above is indeed tractable and all available liquidity is aggregated (Section 4). Finally, we present a design of a fixed rate lending protocol that combines the lending pool and the yield token AMM to provide quotes of the best available interest rates to lenders and borrowers, effectively acting as a lending aggregator (Section 5). This design can be extended to lending vaults (not only pools) managed by curators as well [60, 16]. We provide additional results on the pricing yield tokens for liquid staking/restaking in Appendix [H] and give a preliminary design on how these can give rise to slashing insurance markets. All mathematical proofs are also deferred to the appendices.
@@ 2.1 Model
Definition 2.1 (Yield Bearing Token).
@@
Definition 2.2 (Yield Function).
@@
Definition 2.3 (Yield Tokenizer).
@@
Definition 2.4 (Principal Token).
@@
Definition 2.5 (Yield Token).
@@
Definition 2.6 (Yield Future).
@@ 2.2 Fair pricing
Proof 2.8.
By Ito’s formula, we get
Suppose now, that we have a portfolio consisting of one unit of the yield token $\mathcal{Y}$ and $-\frac{\partial P_{\mathcal{Y}}^{T}}{\partial X^{i}_{t}}$ amounts of the token $\mathcal{A}_{i}$, for all underlying tokens $\{\mathcal{A}_{i}\}$. Let the value of the portfolio in terms of the numeraire $\mathcal{N}$ be $V_{t}$. After a time $dt$, the change in the portfolio value is given by
Substituting Equation [4] in Equation [6] gives us
Also, we know that the value of the investment at time $t$ is given by
$V_{t}$ $=P_{\mathcal{Y}}^{T}-\nabla P_{\mathcal{Y}}^{T}\cdot\bar{X_{t}}$ (10)
where we have $P_{\mathcal{Y}}^{T}(T,x)=0$ as a boundary condition.
To solve this PDE, we first define
[…]
$d\bar{X_{t}}=r(t)\bar{X_{t}}dt+\Sigma_{t}d\bar{W_{t}^{*}}$ (14)
[…]
Taking expectations conditioned on $\bar{X_{t}}=x$ on both sides, using Equation [12] and integrating $s$ from $t$ to $T$, using the fact that the price of the yield token is 00 at maturity, gives us the result.
@@
Although the yield token at a specific maturity $T$ is fungible, it is still distinct from a yield token at some other maturities $T^{\prime}$. The same underlying yield bearing token can give rise to a plethora of yield tokens, each with its own exchanges and liquidity. However, the prices of these seemingly distinct yield tokens are clearly correlated, and one can take advantage of this correlation and unify liquidity across differing maturities. The expression we derived in Theorem 2.7 gives the price of a yield token for a specific maturity $T$. We claim that calculating the price of the yield token across different maturities gives prices that are consistent, in the sense that they do not present a risk-free arbitrage opportunity at any point. More formally, we have the following corollary.
Corollary 2.9.
A portfolio of yield tokens with varying maturities, with the same underlying yield-bearing token, and priced according to Theorem 2.7 cannot have an arbitrage opportunity.
@@
Corollary [2.9] simplifies the task of unifying liquidity across maturities in a yield token market. We can go even further and express any yield token as the sum of yield tokens of smaller maturities in a natural way. For instance, buying a yield token for the time period $[0,T]$ is equivalent to buying $n$ yield tokens for the time periods $[0,t_{1}],[t_{1},t_{2}],\cdots,[t_{n-1},T]$ for any $0<t_{1}<\cdots<t_{n-1}<T$. The yield tokens with the smallest time periods can thus serve as the building blocks to price yield tokens with longer maturities. For instance, the times $t_{1},t_{2},\cdots,t_{n}$ can represent the timestamps of consecutive blocks on the underlying blockchain. However, note that existing protocols do not offer a token that represents the yield in a future time period such as $[t_{1},t_{2}]$. One can write a yield splitting contract, however, that takes a yield token $\mathcal{Y}$ as a deposit and then mints a bunch of yield futures $\{\mathcal{Y}_{i}\}$ for every yield emission in its future. Assuming the underlying yield is emitted at times $t_{1},\cdots,t_{n-1},T$, we define $Y_{i}=\bar{Y}(t_{i},\bar{X_{t_{i}}})\cdot\bar{X_{t_{i}}}$. The random variable $Y_{i}$ represents the yield emitted by the underlying $\mathcal{A}$ at a future time $t_{i}$. By Theorem 2.7, a trader can determine the fair price of a yield future in a straightforward manner.
Corollary 2.10.
@@ 3 Yield risk management
We first zoom in and look at how a trader who wants to express his knowledge can use results from the previous section to make a decision about whether they should go long or short a yield token being traded. In the view of Theorem 2.7, this amounts to having the right model of how interest rate in a lending pool evolves. We give several plausible candidate models that have been developed in DeFi literature recently. Each of these models can be the basis on which Equation [1] can be specified with customized forms for $\bar{\mu}_{t}(.)$ and $\Sigma_{t}(.)$. After that, we zoom out and look at how the welfare of the lending pool as a whole (i.e. the borrowers and lenders) is affected by the presence of the yield tokenizer. Looking at these externalities reveals that risk is allocated in a much better way due to yield tokenization, and it also boosts capital efficiency. We will see how an AMM for a yield token may be designed to provide liquidity in a sustainable manner. This amounts to giving a liquidity provider the right tools to express her knowledge about trader behavior (in particular risk aversion). We also see how a liquid yield token markets give rise to fixed rate lending and money markets.
@@ 3.2 Hedging: a positive externality
Lemma 3.1.
Let $Y$ be the interest amount earned in $\mathcal{N}$ over the period $[0,T]$, as set by the lending pool. Then, under assumptions 1, 2, and 3 the price of the yield token is given by
@@
Proof 3.2.
If the proposed equation does not hold, we have the following cases.
@@
We now move to a more accurate description of the market participants by relaxing Assumption 2. A lender who deposits assets into a pool faces the risk of the pool interest rate dropping below what they expect, while a borrower faces the risk of the interest rate increasing. No matter what methods are used to set interest rates, a lender who has entered the pool agrees on getting paid at least the current equilibrium interest rate, while a borrower agrees to pay at most that amount. If the yield token market is in no arbitrage with the lending pool, we expect risk averse lender to sell, and borrowers to buy yield tokens so as to avoid any changes in the interest rate in the future. We formalize this intuition in the following lemma.
Lemma 3.3.
Let a lender $\mathcal{L}$ deposit $L$ amount of asset and borrower $\mathcal{B}$ borrow $B$ amount of the same asset from a lending pool. Let $Y$ be the interest earned by the lender per unit of asset deposited. They respectively have concave utility functions $U_{\mathcal{L}}$,$U_{\mathcal{B}}$. In the presence of a yield tokenizer, we get the following changes to their behavior.
@@
Proof 3.4.
We have the following two cases.
Taking $\delta_{\mathcal{B}}$,$\delta_{\mathcal{L}}$ as defined above, by Jensen’s inequality, we know that, for any $P_{\mathcal{Y}}^{T}\in[E[Y]-\delta_{\mathcal{L}},E[Y]+\delta_{\mathcal{B}}]$, we have $U_{\mathcal{B}}(-B-B(P_{\mathcal{Y}}^{T})\geq E[U_{\mathcal{B}}(-B-BY/u)]$ and $U_{\mathcal{L}}(W+W(P_{\mathcal{Y}}^{T})\geq E[U_{\mathcal{L}}(L+LY)]$. Thus, we see that a lender prefers to sell her yield tokens, while borrower prefers to buy yield tokens to hedge the interest rate payment on his loan.
We see that risk aversion implies that any lender or borrower is willing to deviate from the fair price to go lower and higher respectively. Thus, if one has the knowledge of how risk averse lenders/borrowers are, one can earn a profit as a market maker by setting prices according to Lemma 3.3. We give recommendations on how this can be done in Section 4, where we discuss how effective hedging can be done when we have limited liquidity in the yield token markets.
@@
Theorem 3.5 tells us that in a completely hedged lending pool, the effects of an interest rate manipulation are not felt by the participating lenders and borrowers. However, the pool still would give a suboptimal view of interest rates to potential lenders/borrowers. In a typical lending pool, the elasticity of loan demand and supply is often quite low [41], and thus there is not a lot of corrective action in response to a rate manipulation. For instance, if the interest rate is increased by an adversary, many potential borrowers would be priced out until lenders put more assets into the pool and drive the interest rate down. However, pairing a lending pool with a yield tokenizer opens up a risk free arbitrage opportunity - namely, a speculator who has a view of what the real equilibrium rate is can lend a small amount of asset, sell the corresponding yield tokens. This would cause the price of yield tokens to drop, thereby forcing underlying interest rate in the pool to drop as well due to the arbitrage opportunities outlined in the proof of Lemma 3.3. Such an arbitrage opportunity would exist as long as there is a significant manipulation by the adversary, thereby bringing the interest rate back to a level which is much closer to attract potential lenders and borrowers alike - the equilibirum value.
@@
Many lending pools present currently have a large gap between borrowing and lending interest rates. So if the yield from such pools is tokenized, borrowers would have to either bid up the prices of yield token in order to achieve complete hedging, or hedge only a part of their loan. This would decrease the welfare of the protocol users as a whole, since risk averse borrowers would only be able to partially hedge their positions. If the borrow absolute interest payment is $\gamma>1$ times the lender’s payment, then we get a worse welfare than complete hedging. More formally, we have the following corollary to Theorem 3.5.
Corollary 3.6.
In the scenario of Theorem 3.5, if the borrowers have to pay a multiplicative premium of $\gamma>1$ on their interest rate, then the welfare of the lending pool participants is non-increasing in $\gamma$.
@@
In previous work [18], the main antidote to adversarial interest rate manipulation has been proposed to be a gap between borrower and lender interest rates. This makes interest rate manipulation less profitable, but also decreases capital efficiency. Thus, an adaptive lending pool has to trade off capital efficiency for interest rate credibility [9]. The presence of a yield tokenizer can help resolve this dilemma. To see this, note that a wider gap makes hedging using yield tokens less effective. This is because a borrower has to buy a yield token at a higher price than the yield being paid as a holder of the token. This would hedge part of the interest rate payment, but would still leave the borrower exposed to some part of the rate manipulation or volatility risk. Thus, if a lending pool is paired with a yield tokenizer, Theorem 3.5 encourages the pool protocol to narrow that gap. It guarantees that doing so would improve user experience (since we have worse user welfare as shown in Corollary [3.6]), while simultaneously protecting the lending pool participants against adversarial manipulation of rates, and volatility.
@@ 4 Designing market makers for yield risk management
In Section 3, we assumed that the market of lenders and borrowers is already in static equilibrium, where all lenders have lent their assets and sold their tokenized position at the same time as the borrowers opening their positions and buying the yield token (Assumption 3). In this section, we analyze what happens when we relax this assumption as well. This is a more realistic scenario - a lender may choose to lend their assets at any time and sell yield tokens, while a borrower may choose to borrow and buy the yield tokens at any other time. There might be other market participants such as speculators or liquidity providers that help make the market. This calls for a market maker that can effectively transfer interest rate risk from the lenders to the borrowers via the yield tokens.
@@
$\psi_{S}(x_{0},y_{0})$ =$=\{(x,y):y=y_{0}+\Delta,x=x_{0}-\Delta p_{\Delta}^{S},\Delta\geq 0\},$ (19)
$\psi_{B}(x_{0},y_{0})$ =$=\{(x,y):y=y_{0}-\Delta,x=x_{0}+\Delta p_{\Delta}^{B},\Delta\geq 0\},$ (20)
@@
A single AMM bonding curve can be thought of as a continuous set of quotes for different trading sizes. A trader interested in selling $\Delta$ amount of the yield token is quoted a price $p_{\Delta}$ based on the changes in reserves permissible under the bonding curve $\psi$. By Theorem 4.3, the bonding curve $\psi$ can be worked out based on the value of yield payments remaining at the time of the trade. Therefore, if the trader decides to sell the yield token just before payment $Y_{j}$, the LP should quote a price based on a bonding curve determined by the distribution of $\sum_{i=j}^{n}Y_{i}$. This implies that the LP should present a menu of $n$ bonding curves, so that the trader can choose when to schedule a buy or sell order. More formally, we can arrive at an optimally efficient menu as follows.
Corollary 4.4.
@@
where the curve $\psi_{j}=\psi(x_{0},y_{0},U_{P},\sum_{j}^{n}Y_{i})$ and the curve function $\psi$ is as defined in Theorem 4.3. Then, each curve $\psi_{j}$ in the menu is optimally efficient.
@@
Given a menu of bonding curves by the liquidity provider, let us now look at the choices of a trader. In particular, assume that the trader wishes to sell an amount $\Delta$ of the yield token. There are $n+1$ choices that he faces - either sell the tokens immediately, or sell them after the first yield payment, or sell them after the first two yield payments,$\cdots$, or sell them after the $n-1^{th}$ payment, or never sell them. Let the prices quoted to the trader by the menu of bonding curves be $p_{1},p_{2},\cdots,p_{n}$ for the first $n$ actions. Corresponding to each of the possible actions, the utility of the trader’s wealth becomes
@@
The results we have shown above give a simple procedure for a prospective liquidity provider to follow. Firstly, the LP brings knowledge of the risk aversion of the lenders/borrower who would be the primary traders the market wants to serve. Secondly, she expresses that knowledge using the constructive definition of an indifference menu as explicated in Lemma 4.6. Thirdly, Theorem 4.8 guarantees that such an indifference menu will also maximize her expected utility over her wealth at maturity. She can then focus on tuning her estimates of how risk averse lenders and borrowers can be as market conditions change, and how traders are reacting to the menu. For instance, she should expect that an indifference menu garners roughly an equal number of trades for each of the $N$ possible actions, if her estimate of risk aversion is accurate. Any net deviation from this (for example, a large number of trades that choose one particular bonding curve in the menu) can be interpreted by her as an indication that she needs to change her estimates of risk aversion for the traders. Many existing data driven market making algorithms can be used for this purpose [61, 62, 14].
@@
Each liquidity provider may express her knowledge of the trader risk aversion differently, by using a different menu of bonding curves, constructed as per Lemma 4.6. What remains to be done is to aggregate these expressions and provide the traders with a unified price quoting mechanism that eases price discovery. This can be done by converting all LP menu bonding curves into demand curves and simply adding them up [58]. This is the effective aggregation strategy used by many state-of-the-art trading interfaces like Tycho [45], Uniswap [77], CoWSwap [70], etc. A result of such aggregation would be that any trader can now view the current price of a yield token and the liquidity around the price (similar to viewing a limit order book) for a menu of buy/sell times into the future, till the maturity of the yield token. This would aid in price discovery of the yield token for lenders and borrowers to hedge their risk, while also assuring them that they are getting the best possible price from the available liquidity. This further enables the creation of fixed rate lending infrastructure, that we outline in the following section.

### v3

@@ 1 Introduction
Although we focus on DeFi lending as the primary case study of where yield tokenization can be used, our results can be extended to other yield-bearing actions such as AMMs and staking as well (Appendix [K]). In automated market makers, liquidity providers (LPs) contribute assets to a pool, earning fees from trades that constitute the yield on their LP tokens. By tokenizing this yield, LPs can sell it to secure a steady cash flow. A more powerful strategy is hedging impermanent loss (loss-vs-holding) [49], by connecting yield tokenization to a demand lending pool [49, 17], allowing LPs to borrow against their LP positions and shift the risk of asset price fluctuation to the lending pool manager. Alternatively, the liquid markets of yield tokens based on LP positions may enable estimation of arbitrage loss (loss-vs-rebalancing) [57], and hedge against uncertainty in the earning of swap fees to compensate for this loss [57]. In liquid staking, tokenizing staking yield similarly allows risk management, though it also introduces some negative externalities (Appendix [K]). A liquid market for such yield tokens enables adversarial validators to profit by shorting the yield and misbehaving in consensus protocols, exacerbating the principal-agent problem that is present there [21]. Nonetheless, yield tokenization enables hedging against other risks, such as Ethereum’s congestion fees: since variable staking rewards depend on transaction/blob fees and validators are known in advance, buyers of yield tokens can hedge against congestion costs while delegators use them to lock in staking rewards [40]. This dual-sided risk transfer resembles the borrower-lender dynamics that we analyze in Section 4 and can be further improved with AMM designs tailored to participants’ risk preferences.
@@
In this work, we make the following key contributions. Firstly, we propose a mathematical model of yield tokenization using stochastic differential equations (Section 2.1). To our knowledge, this is the first formal treatment of yield tokenization and can be used to characterize any yield-bearing token. Secondly, given this model, we derive the fair price of a yield token from a trader’s perspective, in a sense that ensures no arbitrage (Section 2.2). Thirdly, we show how lenders and borrowers in DeFi lending protocols can use yield tokenization to hedge against interest rate uncertainty - thereby improving the welfare of all lending pool participants (Section 3). Fourthly, we design a market maker (using a menu of bonding curves) for yield tokens to be traded efficiently, so that the hedging that we mentioned above is indeed tractable and all available liquidity is aggregated (Section 4). Finally, we present a design of a fixed rate lending protocol that combines the lending pool and the yield token AMM to provide quotes of the best available interest rates to lenders and borrowers, effectively acting as a lending aggregator (Section 5). This design can be extended to lending vaults (not only pools) managed by curators as well [60, 16]. We provide additional results on the pricing yield tokens for liquid staking/restaking in Appendix [K] and give a preliminary design on how these can give rise to slashing insurance markets. All mathematical proofs are also deferred to the appendices.
@@ 2.1 Model
Definition 1 (Yield Bearing Token).
@@
Definition 2 (Yield Function).
@@
Definition 3 (Yield Tokenizer).
@@
Definition 4 (Principal Token).
@@
Definition 5 (Yield Token).
@@
Definition 6 (Yield Future).
@@ 2.2 Fair pricing
Although the yield token at a specific maturity $T$ is fungible, it is still distinct from a yield token at some other maturities $T^{\prime}$. The same underlying yield bearing token can give rise to a plethora of yield tokens, each with its own exchanges and liquidity. However, the prices of these seemingly distinct yield tokens are clearly correlated, and one can take advantage of this correlation and unify liquidity across differing maturities. The expression we derived in Theorem 1 gives the price of a yield token for a specific maturity $T$. We claim that calculating the price of the yield token across different maturities gives prices that are consistent, in the sense that they do not present a risk-free arbitrage opportunity at any point. More formally, we have the following corollary.
Corollary 1.
A portfolio of yield tokens with varying maturities, with the same underlying yield-bearing token, and priced according to Theorem 1 cannot have an arbitrage opportunity.
@@
Corollary [1] simplifies the task of unifying liquidity across maturities in a yield token market. We can go even further and express any yield token as the sum of yield tokens of smaller maturities in a natural way. For instance, buying a yield token for the time period $[0,T]$ is equivalent to buying $n$ yield tokens for the time periods $0,t_{1}],[t_{1},t_{2}],\cdots,[t_{n-1},T]$ for any $0<t_{1}<\cdots<t_{n-1}<T$. The yield tokens with the smallest time periods can thus serve as the building blocks to price yield tokens with longer maturities. For instance, the times $t_{1},t_{2},\cdots,t_{n}$ can represent the timestamps of consecutive blocks on the underlying blockchain. However, note that existing protocols do not offer a token that represents the yield in a future time period such as $t_{1},t_{2}]$. One can write a yield splitting contract, however, that takes a yield token $\mathcal{Y}$ as a deposit and then mints a bunch of yield futures $\{\mathcal{Y}_{i}\}$ for every yield emission in its future. Assuming the underlying yield is emitted at times $t_{1},\cdots,t_{n-1},T$, we define $Y_{i}=\bar{Y}(t_{i},\bar{X_{t_{i}}})\cdot\bar{X_{t_{i}}}$. The random variable $Y_{i}$ represents the yield emitted by the underlying $\mathcal{A}$ at a future time $t_{i}$. By Theorem 1, a trader can determine the fair price of a yield future in a straightforward manner.
Corollary 2.
@@ 3 Yield risk management
We first zoom in and look at how a trader who wants to express his knowledge can use results from the previous section to make a decision about whether they should go long or short a yield token being traded. In the view of Theorem 1, this amounts to having the right model of how interest rate in a lending pool evolves. We give several plausible candidate models that have been developed in DeFi literature recently. Each of these models can be the basis on which Equation [1] can be specified with customized forms for $\bar{\mu}_{t}(.)$ and $\Sigma_{t}(.)$. After that, we zoom out and look at how the welfare of the lending pool as a whole (i.e. the borrowers and lenders) is affected by the presence of the yield tokenizer. Looking at these externalities reveals that risk is allocated in a much better way due to yield tokenization, and it also boosts capital efficiency. We will see how an AMM for a yield token may be designed to provide liquidity in a sustainable manner. This amounts to giving a liquidity provider the right tools to express her knowledge about trader behavior (in particular risk aversion). We also see how a liquid yield token markets give rise to fixed rate lending and money markets.
@@ 3.2 Hedging: a positive externality
Lemma 1.
Let $Y$ be the interest amount earned in $\mathcal{N}$ over the period $[0,T]$, as set by the lending pool. Then, under assumptions Assumption [1], Assumption [2], and Assumption [3] the price of the yield token is given by
@@
We now move to a more accurate description of the market participants by relaxing Assumption Assumption [2]. A lender who deposits assets into a pool faces the risk of the pool interest rate dropping below what they expect, while a borrower faces the risk of the interest rate increasing. No matter what methods are used to set interest rates, a lender who has entered the pool agrees on getting paid at least the current equilibrium interest rate, while a borrower agrees to pay at most that amount. If the yield token market is in no arbitrage with the lending pool, we expect risk averse lender to sell, and borrowers to buy yield tokens so as to avoid any changes in the interest rate in the future. We formalize this intuition in the following lemma.
Lemma 2.
Let a lender $\mathcal{L}$ deposit $L$ amount of asset and borrower $\mathcal{B}$ borrow $B$ amount of the same asset from a lending pool. Let $Y$ be the interest earned by the lender per unit of asset deposited. They respectively have concave utility functions $_{\mathcal{L}},UU_{\mathcal{B}}$. In the presence of a yield tokenizer, we get the following changes to their behavior.
@@
We see that risk aversion implies that any lender or borrower is willing to deviate from the fair price to go lower and higher respectively. Thus, if one has the knowledge of how risk averse lenders/borrowers are, one can earn a profit as a market maker by setting prices according to Lemma [2]. We give recommendations on how this can be done in Section 4, where we discuss how effective hedging can be done when we have limited liquidity in the yield token markets.
@@
Theorem 2 tells us that in a completely hedged lending pool, the effects of an interest rate manipulation are not felt by the participating lenders and borrowers. However, the pool still would give a suboptimal view of interest rates to potential lenders/borrowers. In a typical lending pool, the elasticity of loan demand and supply is often quite low [41], and thus there is not a lot of corrective action in response to a rate manipulation. For instance, if the interest rate is increased by an adversary, many potential borrowers would be priced out until lenders put more assets into the pool and drive the interest rate down. However, pairing a lending pool with a yield tokenizer opens up a risk free arbitrage opportunity - namely, a speculator who has a view of what the real equilibrium rate is can lend a small amount of asset, sell the corresponding yield tokens. This would cause the price of yield tokens to drop, thereby forcing underlying interest rate in the pool to drop as well due to the arbitrage opportunities outlined in the proof of Lemma [2]. Such an arbitrage opportunity would exist as long as there is a significant manipulation by the adversary, thereby bringing the interest rate back to a level which is much closer to attract potential lenders and borrowers alike - the equilibirum value.
@@
Many lending pools present currently have a large gap between borrowing and lending interest rates. So if the yield from such pools is tokenized, borrowers would have to either bid up the prices of yield token in order to achieve complete hedging, or hedge only a part of their loan. This would decrease the welfare of the protocol users as a whole, since risk averse borrowers would only be able to partially hedge their positions. If the borrow absolute interest payment is $\gamma>1$ times the lender’s payment, then we get a worse welfare than complete hedging. More formally, we have the following corollary to Theorem 2.
Corollary 3.
In the scenario of Theorem 2, if the borrowers have to pay a multiplicative premium of $\gamma>1$ on their interest rate, then the welfare of the lending pool participants is non-increasing in $\gamma$.
@@
In previous work [18], the main antidote to adversarial interest rate manipulation has been proposed to be a gap between borrower and lender interest rates. This makes interest rate manipulation less profitable, but also decreases capital efficiency. Thus, an adaptive lending pool has to trade off capital efficiency for interest rate credibility [9]. The presence of a yield tokenizer can help resolve this dilemma. To see this, note that a wider gap makes hedging using yield tokens less effective. This is because a borrower has to buy a yield token at a higher price than the yield being paid as a holder of the token. This would hedge part of the interest rate payment, but would still leave the borrower exposed to some part of the rate manipulation or volatility risk. Thus, if a lending pool is paired with a yield tokenizer, Theorem 2 encourages the pool protocol to narrow that gap. It guarantees that doing so would improve user experience (since we have worse user welfare as shown in Corollary [3]), while simultaneously protecting the lending pool participants against adversarial manipulation of rates, and volatility.
@@ 4 Designing market makers for yield risk management
In Section 3, we assumed that the market of lenders and borrowers is already in static equilibrium, where all lenders have lent their assets and sold their tokenized position at the same time as the borrowers opening their positions and buying the yield token (Assumption Assumption [3]). In this section, we analyze what happens when we relax this assumption as well. This is a more realistic scenario - a lender may choose to lend their assets at any time and sell yield tokens, while a borrower may choose to borrow and buy the yield tokens at any other time. There might be other market participants such as speculators or liquidity providers that help make the market. This calls for a market maker that can effectively transfer interest rate risk from the lenders to the borrowers via the yield tokens.
@@
$\psi_{S}(x_{0},y_{0})$ $=\{(x,y):y=y_{0}+\Delta,x=x_{0}-\Delta p_{\Delta}^{S},\Delta\geq 0\},$ (6)
$\psi_{B}(x_{0},y_{0})$ $=\{(x,y):y=y_{0}-\Delta,x=x_{0}+\Delta p_{\Delta}^{B},\Delta\geq 0\},$ (7)
@@
A single AMM bonding curve can be thought of as a continuous set of quotes for different trading sizes. A trader interested in selling $\Delta$ amount of the yield token is quoted a price $p_{\Delta}$ based on the changes in reserves permissible under the bonding curve $\psi$. By Theorem 3, the bonding curve $\psi$ can be worked out based on the value of yield payments remaining at the time of the trade. Therefore, if the trader decides to sell the yield token just before payment $Y_{j}$, the LP should quote a price based on a bonding curve determined by the distribution of $\sum_{i=j}^{n}Y_{i}$. This implies that the LP should present a menu of $n$ bonding curves, so that the trader can choose when to schedule a buy or sell order. More formally, we can arrive at an optimally efficient menu as follows.
Corollary 4.
@@
where the curve $\psi_{j}=\psi(x_{0},y_{0},U_{P},\sum_{j}^{n}Y_{i})$ and the curve function $\psi$ is as defined in Theorem 3. Then, each curve $\psi_{j}$ in the menu is optimally efficient.
@@
[…]
@@
The results we have shown above give a simple procedure for a prospective liquidity provider to follow. Firstly, the LP brings knowledge of the risk aversion of the lenders/borrower who would be the primary traders the market wants to serve. Secondly, she expresses that knowledge using the constructive definition of an indifference menu as explicated in Lemma [5]. Thirdly, Theorem 4 guarantees that such an indifference menu will also maximize her expected utility over her wealth at maturity. She can then focus on tuning her estimates of how risk averse lenders and borrowers can be as market conditions change, and how traders are reacting to the menu. For instance, she should expect that an indifference menu garners roughly an equal number of trades for each of the $N$ possible actions, if her estimate of risk aversion is accurate. Any net deviation from this (for example, a large number of trades that choose one particular bonding curve in the menu) can be interpreted by her as an indication that she needs to change her estimates of risk aversion for the traders. Many existing data driven market making algorithms can be used for this purpose [61, 62, 14].
@@
Each liquidity provider may express her knowledge of the trader risk aversion differently, by using a different menu of bonding curves, constructed as per Lemma [5]. What remains to be done is to aggregate these expressions and provide the traders with a unified price quoting mechanism that eases price discovery. This can be done by converting all LP menu bonding curves into demand curves and simply adding them up [58]. This is the effective aggregation strategy used by many state-of-the-art trading interfaces like Tycho [45], Uniswap [77], CoWSwap [70], etc. A result of such aggregation would be that any trader can now view the current price of a yield token and the liquidity around the price (similar to viewing a limit order book) for a menu of buy/sell times into the future, till the maturity of the yield token. This would aid in price discovery of the yield token for lenders and borrowers to hedge their risk, while also assuring them that they are getting the best possible price from the available liquidity. This further enables the creation of fixed rate lending infrastructure, that we outline in the following section.